# BOOM: Benchmarking Out-Of-distribution Molecular Property Predictions of Machine Learning Models

### v1

@@ Abstract
Advances in deep learning and generative modeling have driven interest in data-driven molecule discovery pipelines, whereby machine learning (ML) models are used to filter and design novel molecules without requiring prohibitively expensive first-principles simulations. Although the discovery of novel molecules that extend the boundaries of known chemistry requires accurate out-of-distribution (OOD) predictions, ML models often struggle to generalize OOD. Furthermore, there are currently no systematic benchmarks for molecular OOD prediction tasks. We present BOOM, benchmarks for out-of-distribution molecular property predictions—a benchmark study of property-based out-of-distribution models for common molecular property prediction models. We evaluate more than 140 combinations of models and property prediction tasks to benchmark deep learning models on their OOD performance. Overall, we do not find any existing models that achieve strong OOD generalization across all tasks: even the top performing model exhibited an average OOD error 3x larger than in-distribution. We find that deep learning models with high inductive bias can perform well on OOD tasks with simple, specific properties. Although chemical foundation models with transfer and in-context learning offer a promising solution for limited training data scenarios, we find that current foundation models do not show strong OOD extrapolation capabilities. We perform extensive ablation experiments to highlight how OOD performance is impacted by data generation, pre-training, hyperparameter optimization, model architecture, and molecular representation. We propose that developing ML models with strong OOD generalization is a new frontier challenge in chemical ML model development. This open-source benchmark will be made available on Github.
@@ 1 Introduction
Molecular discovery is the process by which novel molecular structures with desirable application-specific properties are identified. Given the immense total search space of hypothetical small molecules enumerating approximately $10^{60}$ hypothetical molecules and 100 billion enumerated molecules, molecule discovery has increasingly relied upon machine learning models to efficiently navigate this space.[1, 2, 3] Typically, molecule discovery is performed by first training a machine learning (ML) model on a molecule property dataset to learn the property-structure relationship. Then, this trained ML model is used to discover new molecules either by screening a list of enumerated molecules or by guiding a generative model towards molecules of interest [4].
@@
Molecule discovery is inherently an out-of-distribution (OOD) prediction problem. For the discovered molecules to constitute an exciting chemical discovery, the molecules need to either i) exhibit properties that extrapolate beyond those of the known molecules in the training dataset, or ii) possess a new chemical substructure that was previously not considered for the application of interest. In either case, the success of the molecule discovery campaign is dependent on the machine learning model’s ability to make accurate predictions on samples that do not follow the same distribution as the known molecules (training data).
@@
Despite the importance of OOD performance to the problem of molecule discovery, the OOD performance of commonly used ML models for molecular property prediction has yet to be systematically explored. The majority of the standardized benchmarks used to assess the performance of chemical property prediction models do not include evaluations of model performance in the case where the test set is drawn from a different distribution than the training data. As a result of the lack of OOD chemistry benchmarks, the development of chemistry ML models are currently driven primarily by maximizing in-distribution performance, which may be hurting model generalization. The lack of OOD chemistry benchmarks has also hindered our understanding of how to develop generalizable chemistry foundation models. Currently, there is little empirical knowledge about how choices regarding the pretraining task, model architecture, and/or dataset diversity impact the generalization performance of chemistry foundation models that are expected to generalize across all chemical systems.
@@
In this work, we develop BOOM, benchmarks for out-of-distribution molecular property predictions, a standardized benchmark for assessing the OOD generalization performance of molecule property prediction models.
1.1 Main Findings
Our work consists of the following main contributions:
@@ 1.1 Main Findings
  • We develop a robust methodology for evaluating the performance of chemical property prediction models to extrapolate to property values beyond their training distribution. Notably, this methodology is developed in a general manner, allowing it to apply to any material property dataset regardless of the specific model architecture, material property, or chemical system.
@@
  • We perform the first large-scale benchmarking of the OOD performance of state-of-the-art ML chemical property prediction models. Across 10 diverse OOD tasks and 12 ML models, we do not find any existing models that show strong OOD generalization across all tasks. We therefore put forth BOOM OOD property prediction as a frontier challenge for chemical foundation models.
@@
  • Our work highlights insights into how pretraining strategies, model architecture, molecule representation, and data augmentation impact OOD performance. These findings point towards strategies for the chemistry community to achieve chemical foundation models with strong OOD generalization across all chemical systems.
2 BOOM Overview
In general, one can define OOD with respect to either the model inputs (holding out a region of chemical space as the OOD test split) or with respect to the model outputs (holding out a range of chemical property values). In this work, we adopt the latter approach of benchmarking the performance of the models to extrapolate to property values not seen in training. Following the OOD definitions outlined by Farquhar et al., we here define OOD as a complement distribution with respect to the targets [5, 6]. Specifically, given a molecule property dataset of chemical structures and their numerical property values, we create our OOD test set to consist of numerical values on the tail ends of the numerical property distribution (see Figure 1). In this way, our OOD benchmarking is directly aligned with the molecule discovery task in that it allows us to evaluate the consistency of ML models to discover molecules with state-of-the-art properties that extrapolate beyond the training data.
2.1 Datasets
Overall, BOOM consists of 10 unique molecular property datasets. We collect 8 molecular property datasets from the QM9 Dataset: isotropic polarizability ($\alpha$), heat capacity ($C_{v}$), HOMO energy, LUMO energy, HOMO-LUMO gap, dipole moment ($\mu$) electronic spatial extent ($R^{2}$) and zero point vibrational energy (zpve).[7] Each property in the QM9 dataset was determined from Density Functional Theory (DFT) calculations on the same set of 133,886 small molecules made up of CHONF atoms. We also benchmark on the 10k Dataset, which consists of two molecular properties calculated by DFT: density and solid heat of formation.[8] The 10k Dataset was sourced from 10,206 experimentally synthesized CHON small molecules from the Cambridge Crystal Structure Dataset.
2.2 OOD Splitting
We generate our training, ID, and OOD splits based on the property distribution. For each of the 10 molecular properties, we generate OOD splits by first fitting a kernel density estimator (with Gaussian kernel) to the property values and obtain the probability of a molecule given its property. We select the molecules with the lowest probabilities for the OOD split for that property. This results in selecting the molecules at the tail end of the distribution for typical molecular property distributions since these molecules will have low densities in property space. Unlike partitioning by cut-off values, this method of splitting allows us to capture low-probability samples for general distributions that aren’t necessarily unimodal.
For QM9 we take the lowest 10% of the probability scores as predicted by the kernel density estimator for the OOD set. We take the lowest 1000 molecules for the 10K dataset. We then randomly sample molecules from the remaining molecules to generate the ID test set. We sample 10% of the molecules in the case of QM9 and 5% of the molecules for ID test split for 10K data. The remaining molecules are used for training and fine-tuning. Two example OOD splits are provided in Figure 1, whereas the remaining datasets can be found in the Appendix 8.1.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Molecular discovery pipelines have increasingly relied upon machine learning (ML) models (bohacek_art_1996; reymond_chemical_2015; kailkhura2019reliable). These models discover new molecules by either screening a list of enumerated molecules or by guiding a generative model towards molecules of interest (wang2023scientific). Molecular discovery is inherently an out-of-distribution (OOD) prediction problem, since the molecules need to either (i) exhibit properties that extrapolate beyond the training dataset, or (ii) possess a previously unconsidered chemical substructure. In either case, success depends on the learned model’s ability to make accurate predictions on samples that are not in the same distribution as the training data.
@@
Despite the importance of OOD performance to real-world molecular discovery, the OOD performance of common ML models for molecular property prediction has yet to be systematically explored. Due to the lack of standardized splits for testing models, especially splits based on the data distribution, we believe that current ML models are optimizing in-distribution performance on insufficiently challenging datasets that do not adequately measure real-world performance. Currently, little empirical knowledge exists about how choices regarding the pretraining task, model architecture, and/or dataset diversity impact the generalization performance of chemistry foundation models that are expected to generalize across all chemical systems.
@@
In this work, we develop BOOM, benchmarks for out-of-distribution molecular property predictions, a standardized benchmark for assessing the OOD generalization performance of molecule property prediction models. Our work consists of the following main contributions:
@@
  • We develop a general and robust methodology for evaluating the performance of chemical property prediction models for property values beyond their training distribution. We introduce OOD-specific metrics such as binned $R^{2}$ to allow comparisons of OOD performance across all models.
@@
  • We perform the first large-scale OOD performance benchmarking of state-of-the-art ML chemical property prediction models. Across 10 diverse OOD tasks and 15 models, we do not find any existing models that show strong OOD generalization across all tasks. We therefore put forth BOOM OOD property prediction as a frontier challenge for chemical foundation models.
@@
  • Our work highlights insights into how pretraining strategies, model architecture, molecular representation, and data augmentation impact OOD performance. These findings point towards strategies for the chemistry community to achieve chemical foundation models with strong OOD generalization across all chemical systems.
2 BOOM
Defining Out-of-distribution. Consider a supervised dataset $\mathcal{D}$ with $N$ molecules $\mathcal{M}\in\{\mathcal{M}_{1},\mathcal{M}_{2},...,\mathcal{M}_{N}\}$ and associated labels or properties $y\in\{y_{1},y_{2},...,y_{N}\}$. The problem of out-of-distribution prediction can be defined as the mismatch in the probability distribution, $P$ of the training and test sets, $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ such that,
$P(\mathcal{M},y|\mathcal{D}_{test})\neq P(\mathcal{M},y|\mathcal{D}_{train})$ (1)
The key question is defining the density function $P(\mathcal{M},y)$ over a set of molecules and their respective properties. The density can be defined over the chemical structure or molecule features, or over the properties. Formally, we define out-of-distribution as low-density regions over the property space, such that:
$0<P(y_{\text{test}})\leq P(y_{\text{train}})$ (2)
farquhar2022out define this as a complement distribution conditioned on the targets. This is known as concept or label shift as well (liu2024efficacy). While we focus on designing splits with a concept shift, it is important to note that depending on the property, this may result in a covariate shit, resulting in a structural or chemical imbalance. The probability density over the labels is determined using kernel density estimation (KDE), allowing us to generalize to multimodal distributions. The split strategy algorithm for each dataset is detailed in Appendix A.1. The lowest probability samples from the KDE estimated distribution are held-out (see Fig. 1) to evaluate the consistency of ML models to discover molecules with state-of-the-art properties that extrapolate beyond the training data.
Datasets. BOOM consists of 10 quantum chemical molecular property datasets derived from QM9 (ramakrishnan_QM9_2014) and the 10k Dataset (antoniuk_active_2025), derived from the Cambridge Structural Database. The 10k Dataset was sourced from 10,206 experimentally synthesized, small organic molecules and contains the density functional theory calculated values of their molecular density and solid heat of formation (HoF). We collect 8 molecular property datasets from the QM9 Dataset: isotropic polarizability ($\alpha$), heat capacity ($C_{v}$), highest occupied molecular orbital (HOMO) energy, lowest unoccupied molecular orbital (LUMO) energy, HOMO-LUMO gap, dipole moment ($\mu$), electronic spatial extent ($\left<R^{2}\right$>), and zero point vibrational energy (ZPVE). We also select a random subset of the dataset to serve as the ID test set, detailed in Appendix A. To further expand the application space of BOOM, we also perform benchmarking on the Lipophilicity dataset(moleculenet) of 4,200 experimental measurements of the octanol/water distribution coefficient, which is of relevance for drug compounds. The inclusion of the Lipophilicity dataset serves as an exemplary dataset for performing OOD evaluations on experimentally measured properties, rather than only computed physicochemical properties (See Table 9).